# Spin precession effects in the phasing formula of eccentric compact binary inspirals till the second post-Newtonian order

### v1

@@ II.4 Resummation of the TaylorT2 phase
$-\frac{46697\nu}{37696}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{226079}{5126656}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+\frac{10479\nu}{640832}+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}$
@@
$-\frac{10479}{1281664}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{3115}{4352}-\frac{3115\nu}{2176}\biggr)\kappa_{A}-\frac{10479\delta}{640832}+\frac{3115\delta\kappa_{S}}{4352}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+\frac{10479\nu}{640832}$
@@
$+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{10479}{1281664}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}+\frac{1435105\nu}{349184}+\biggl(\frac{18015\nu}{2816}$
@@
$-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{18015\nu}{1408}-\frac{18015}{2816}\biggr)\kappa_{A}+\frac{73355\delta}{349184}-\frac{18015\delta\kappa_{S}}{2816}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}-\frac{1581815\nu}{349184}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}$
@@
$+\frac{54135\nu}{349184}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{6005}{2816}-\frac{6005\nu}{1408}\biggr)\kappa_{A}-\frac{54135\delta}{349184}+\frac{6005\delta\kappa_{S}}{2816}\biggr]$
@@
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}+\frac{54135\nu}{349184}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl[\biggl(\frac{215775}{31966208}$
$-\frac{215775\nu}{15983104}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{215775\delta}{15983104}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{215775\nu}{15983104}-\frac{215775}{31966208}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$-\frac{215775\delta}{15983104}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{215775}{31966208}-\frac{215775\nu}{15983104}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{215775\nu}{15983104}-\frac{215775}{31966208}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{\delta\kappa_{A}}{4}+\frac{531\nu}{19840}+\biggl(\frac{1}{4}-\frac{\nu}{2}\biggr)\kappa_{S}-\frac{531}{39680}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{1}{2}-\nu\biggr)\kappa_{A}-\frac{531\delta}{19840}+\frac{\delta\kappa_{S}}{2}\biggr]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{\delta\kappa_{A}}{4}+\frac{531\nu}{19840}+\biggl(\frac{1}{4}-\frac{\nu}{2}\biggr)\kappa_{S}-\frac{531}{39680}\biggr]$ (22f)
@@ II.5 Spin effects in the Fourier domain phase (SUA)
From Eq. (108) of [77] it is clear that the SUA Fourier domain phasing formula remains unchanged compared to SPA. Owing to this simplification, we find the Fourier domain phasing formula to be
@@
where
$\Psi_{\rm NS,circ}$ $=1+y^{2}\left(\frac{3715}{756}+\frac{55\nu}{9}\right)-16\pi y^{3}+y^{4}\left(\frac{15293365}{508032}+\frac{27145\nu}{504}+\frac{3085\nu^{2}}{72}\right)\,$, (25a)
$\Psi_{\rm SO,circ}$ $=\frac{113\delta}{3}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\left(\frac{113}{3}-\frac{76\nu}{3}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\,$, (25b)
$\Psi_{\rm SS,circ}$ $=\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-75\delta\kappa_{A}+\frac{96495\nu}{992}+(150\nu-75)\kappa_{S}+\frac{1465}{1984}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-75\delta\kappa_{A}-\frac{99425\nu}{992}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl(\frac{1625}{181288}-\frac{1625\nu}{90644}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{1625\delta}{90644}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{1625\nu}{90644}$
$-\frac{1625}{181288}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}-\frac{1625\delta}{90644}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{1625}{181288}-\frac{1625\nu}{90644}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{1625\nu}{90644}$
$-\frac{1625}{181288}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[25\delta\kappa_{A}+\frac{2705\nu}{992}+(25-50\nu)\kappa_{S}-\frac{2705}{1984}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(50-100\nu)\kappa_{A}-\frac{2705\delta}{992}+50\delta\kappa_{S}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[25\delta\kappa_{A}+\frac{2705\nu}{992}+(25-50\nu)\kappa_{S}-\frac{2705}{1984}\biggr]\,$, (25c)
$\Psi_{\rm NS,ecc}$ $=\frac{650}{731}\left(\frac{y_{0}}{y}\right)^{19/3}+y^{2}\left[\left(\frac{y_{0}}{y}\right)^{25/3}\left(\frac{920725}{368424}-\frac{64025\nu}{13158}\right)+\left(\frac{y_{0}}{y}\right)^{19/3}\left(-\frac{5776985}{1044288}+\frac{467015\nu}{37296}\right)\right]$
$+y^{3}\left[-\frac{79691\pi}{6120}\left(\frac{y_{0}}{y}\right)^{19/3}+\frac{122525\pi}{26316}\left(\frac{y_{0}}{y}\right)^{28/3}\right]+y^{4}\biggl[\left(\frac{y_{0}}{y}\right)^{25/3}\left(-\frac{16366198505}{1052642304}+\frac{205093295\nu}{3132864}\right$.
$\left.-\frac{92001955\nu^{2}}{1342656}\right)+\left(\frac{y_{0}}{y}\right)^{31/3}\left(-\frac{387806575}{1114114176}-\frac{1267825\nu}{195048}+\frac{5900375\nu^{2}}{473688}\right)+\left(\frac{y_{0}}{y}\right)^{19/3}\left(\frac{14983962085}{2078866944}\right$.
$\left.-\frac{12626215\nu}{220968}+\frac{66423155\nu^{2}}{883872}\right)\biggr]\,$, (25d)
$\Psi_{\rm SO,ecc}$ $=\left(\frac{y_{0}}{y}\right)^{19/3}\left[\frac{45793\delta}{1836}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\left(\frac{45793}{1836}-\frac{29659\nu}{2295}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right]+\left(\frac{y_{0}}{y}\right)^{28/3}\left[-\frac{51025\delta}{19737}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right$.
$\left.+\left(-\frac{51025}{19737}+\frac{35750\nu}{19737}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right]\,$, (25e)
$\Psi_{\rm SS,ecc}$ $=\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[\frac{28925\delta\kappa_{A}}{11696}-\frac{1086475\nu}{333336}+\biggl(\frac{28925}{11696}-\frac{28925\nu}{5848}\biggr)\kappa_{S}+\frac{4225}{41667}\biggr]$
@@
$+\frac{951275\nu}{333336}+\biggl(\frac{28925}{11696}-\frac{28925\nu}{5848}\biggr)\kappa_{S}+\frac{4225}{41667}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[-\frac{28925\delta\kappa_{A}}{35088}-\frac{4225\nu}{111112}+\biggl(\frac{28925\nu}{17544}$
@@
$-\frac{28925}{35088}\biggr)\kappa_{S}+\frac{4225}{222224}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{28925\nu}{8772}-\frac{28925}{17544}\biggr)\kappa_{A}+\frac{4225\delta}{111112}-\frac{28925\delta\kappa_{S}}{17544}\biggr]$
@@
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[-\frac{28925\delta\kappa_{A}}{35088}-\frac{4225\nu}{111112}+\biggl(\frac{28925\nu}{17544}-\frac{28925}{35088}\biggr)\kappa_{S}+\frac{4225}{222224}\biggr]\biggr\}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{294955\delta\kappa_{A}}{10912}+\frac{3067286795\nu}{87274176}+\biggl(\frac{294955\nu}{5456}-\frac{294955}{10912}\biggr)\kappa_{S}+\frac{35057425}{174548352}\biggr]$

### v2

@@ II.4 Resummation of the TaylorT2 phase
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1918567\nu}{2563328}}+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{10479}{1281664}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{3115}{4352}-\frac{3115\nu}{2176}\biggr)\kappa_{A}-\frac{10479\delta}{640832}+\frac{3115\delta\kappa_{S}}{4352}\biggr]$
@@
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{3115\nu}{4352}}+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{10479}{1281664}\biggr]\biggr\}$
@@
$++\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{28925\nu}{15872}}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{18015\nu}{1408}-\frac{18015}{2816}\biggr)\kappa_{A}+\frac{73355\delta}{349184}-\frac{18015\delta\kappa_{S}}{2816}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{6315\nu}{2816}}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{426445\nu}{174592}}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{6005}{2816}-\frac{6005\nu}{1408}\biggr)\kappa_{A}-\frac{54135\delta}{349184}+\frac{6005\delta\kappa_{S}}{2816}\biggr]$
@@
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{6005\nu}{2816}}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]\biggr\}$
$+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl[\biggl(\frac{215775}{31966208}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{215775\nu}{7991552}}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{215775\delta}{15983104}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)$
$+\biggl({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{215775\nu}{7991552}}-\frac{215775}{31966208}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}-\frac{215775\delta}{15983104}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{215775}{31966208}}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{215775}{31966208}}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{\delta\kappa_{A}}{4}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{5491\nu}{9920}}+\biggl(\frac{1}{4}-\frac{\nu}{2}\biggr)\kappa_{S}-\frac{531}{39680}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{1}{2}-\nu\biggr)\kappa_{A}-\frac{531\delta}{19840}+\frac{\delta\kappa_{S}}{2}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{\delta\kappa_{A}}{4}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{\nu}{2}}+\biggl(\frac{1}{4}-\frac{\nu}{2}\biggr)\kappa_{S}-\frac{531}{39680}\biggr]$ (23f)
Note that these coefficients up to $\mathcal{O}(e_{0}^{8})$ are given in [104].
@@ II.5 Spin effects in the Fourier domain phase (SUA)
From Eq. (108) of [87] it is clear that the SUA Fourier domain phasing formula remains unchanged compared to the stationary phase approximation (SPA). Owing to this simplification, as was shown in Sec. VI E of Ref. [53], we find the Fourier domain phasing formula to be
@@
where $\Psi_{\rm NS,circ}$ and $\Psi_{\rm NS,ecc}$ are provided in [105], $\Psi_{\rm SO,circ}$ in [70], and $\Psi_{\rm SO,ecc}$ in [101]. Rest of the coefficients are as follows-
$\Psi_{\rm SS,circ}$ $=\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-75\delta\kappa_{A}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{22095\nu}{496}}+(150\nu-75)\kappa_{S}+\frac{1465}{1984}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-75\delta\kappa_{A}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{95\nu}{2}}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl(\frac{1625}{181288}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1625\nu}{45322}}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{1625\delta}{90644}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1625\nu}{45322}}$
$-\frac{1625}{181288}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}-\frac{1625\delta}{90644}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1625}{181288}}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}$
$-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1625}{181288}}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[25\delta\kappa_{A}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{27505\nu}{496}}+(25-50\nu)\kappa_{S}-\frac{2705}{1984}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(50-100\nu)\kappa_{A}-\frac{2705\delta}{992}+50\delta\kappa_{S}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[25\delta\kappa_{A}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-50\nu}+(25-50\nu)\kappa_{S}-\frac{2705}{1984}\biggr]\,$, (26a)
$\Psi_{\rm SS,ecc}$ $=\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[\frac{28925\delta\kappa_{A}}{11696}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{524225\nu}{333336}}+\biggl(\frac{28925}{11696}-\frac{28925\nu}{5848}\biggr)\kappa_{S}+\frac{4225}{41667}\biggr]$
@@
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{6825\nu}{5848}}+\biggl(\frac{28925}{11696}-\frac{28925\nu}{5848}\biggr)\kappa_{S}+\frac{4225}{41667}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[-\frac{28925\delta\kappa_{A}}{35088}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{574925\nu}{333336}}$
$+\biggl(\frac{28925\nu}{17544}-\frac{28925}{35088}\biggr)\kappa_{S}+\frac{4225}{222224}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{28925\nu}{8772}-\frac{28925}{17544}\biggr)\kappa_{A}+\frac{4225\delta}{111112}-\frac{28925\delta\kappa_{S}}{17544}\biggr]$
@@
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[-\frac{28925\delta\kappa_{A}}{35088}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\frac{28925\nu}{17544}}+\biggl(\frac{28925\nu}{17544}-\frac{28925}{35088}\biggr)\kappa_{S}+\frac{4225}{222224}\biggr]\biggr\}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{294955\delta\kappa_{A}}{10912}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{64385155\nu}{3967008}}+\biggl(\frac{294955\nu}{5456}-\frac{294955}{10912}\biggr)\kappa_{S}$
$+\frac{35057425}{174548352}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{294955\nu}{2728}-\frac{294955}{5456}\biggr)\kappa_{A}+\frac{35057425\delta}{87274176}-\frac{294955\delta\kappa_{S}}{5456}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{294955\delta\kappa_{A}}{10912}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{92935\nu}{5456}}+\biggl(\frac{294955\nu}{5456}-\frac{294955}{10912}\biggr)\kappa_{S}+\frac{35057425}{174548352}\biggr]$